COMPUTATIONAL MOLECULAR SPECTROSCOPY OF FeCN IN THE ^{6}Δ ELECTRONIC GROUND STATE

\underline{TSUNEO HIRANO}, REI OKUDA, and UMPEI NAGASHIMA, Research Institute for Computational Sciences, National Institute of Advanced Industrial Science and Technology, 1-1-1 Umezono, Tsukuba, Ibaraki 305-8568, Japan; MICHIKO AMANO, YUKARI MITSUI, SACHIKO S. ITONO, Department of Chemistry, Faculty of Science, Ochanomizu University, 2-1-1 Otsuka, Bunkyo-ku, Tokyo 112-8610, Japan; PER JENSEN, Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

We have previously reported a computational molecular spectroscopic study of ^{6}Δ_{i} FeNC,^{a} where we showed that the experimentally derived, too-short C-N bond length (r_{e}(C-N) = 1.03(8) Å)^{b} can be ascribed to an inadequate treatment of the large amplitude bending motion in the experimental determination of r_{0}. Here, we report analogous calculations for ^{6}Δ_{i} FeCN.^{c} Based on the three-dimensional potential energy surface calculated at the MR-SDCI+Q+E_{rel}/[Roos ANO (Fe), aug-cc-pVQZ (C, N)] level of theory, the standard spectroscopic parameters of Fe^{12}CN and Fe^{13}CN are derived by perturbation methods, and ro-vibrationally averaged bond lengths ⟨r⟩ have been predicted as expectation values obtained with rovibrational wavefunctions from the MORBID program. Some of the spectroscopic constants thus determined are: r_{e}(Fe-C) = 2.048 Å and r_{e}(C-N) = 1.168 Å, ω_{1} = 2179 cm^{−1}, ω_{2} = 173 cm^{−1}, ω_{3} = 420 cm^{−1}, dipole moment = 4.59 D, spin-orbit coupling constant A_{SO} = −83 cm^{−1}, ⟨r(Fe-C)⟩_{0} = 2.082 Å, and ⟨r(C-N)⟩_{0} = 1.172 Å. In variational MORBID calculations, rovibronic energy levels are determined, and some vibrational bands are simulated. The bending potential is shallow, and the MORBID calculations show that the zero-point averaged structure is bent with the expectation value ⟨^{}(Fe-C-N)⟩_{0} = 170(5)^{∘} (where the number in parentheses is the quantum-mechanical uncertainty). We compare the ^{6}Δ_{i} FeCN results with those obtained for ^{6}Δ_{i} FeNC. Since there are no experimental spectroscopic data available for FeCN, we hope that the predictions made here may be useful in the experimental investigation of this molecule.

aT. Hirano, R. Okuda, U. Nagashima, V. Špirko, and P. Jensen, J. Mol. Spectrosc., 236, 234-247 (2006).

bJ. Lie and P. J. Dagdigian, J. Chem. Phys., 114, 2137-2143 (2001).

cT. Hirano, M. Amano, Y. Mitsui, S.S. Itono, R. Okuda, U. Nagashima, and P. Jensen, J. Mol. Spectrosc., in press (2007).